Problem 3-38 (Part Level Submission) Cullumber Company produces a molded briefcase that is distributed to luggage stores. The following operating data for the current year has been accumulated for planning purposes. Sales price $41.00 Variable cost of goods sold 13.00 Variable selling expenses 11.60 Variable administrative expenses 4.00 Annual fixed expenses Overhead $22,315,800 Selling expenses 4,434,550 Administrative expenses 9,298,250 Cullumber can produce 4,291,500 cases a year. The projected net income for the coming year is expected to be $5,149,800. Cullumber is subject to a 40% income tax rate. During the planning sessions, Cullumber’s managers have been reviewing costs and expenses. They estimate that the company’s variable cost of goods sold will increase 15% in the coming year and that fixed administrative expenses will increase by $429,150. All other costs and expenses are expected to remain the same. Cullumber Company’s managers are considering expanding the product line by introducing a leather briefcase. The new briefcase is expected to sell for $91.00; variable costs would amount to $37.00 per briefcase. If Cullumber introduces the leather briefcase, the company will incur an additional $858,300 per year in advertising costs. Cullumber’s marketing department has estimated that one new leather briefcase would be sold for every four molded briefcases. Collapse question part (a) If managers decide to introduce the new leather briefcase, how many units of each briefcase would be required to break even in the coming year? Cost of goods sold for the molded briefcase is expected to be $14.95 per unit. (For computational purposes round contribution margin per unit to 2 decimal places, e.g. 0.38. Round answers to 0 decimal places, e.g. 25,000.)
Answer #1

Contribution margin molded briefcases = 41 - 14.95 -11.60 -4 = 10.45 * 4/5 = 8.36

Contribution margin leather briefcases = 91 - 37 = 54 *1/5 = 10.80

Total fixed cost = 22,315,800 + 4,434,550 + 9,298,250 + 429,150 + 858,300 = 37,336,050

Break even sales = 37,336,050 / (8.36+10.80) = 1,948,646

Break even sales  molded briefcases = 1,948,646 * 4/ 5 = 1558917

Break even sales leather briefcases = 1,948,646 * 1/ 5 = 389,729